# Provably Efficient Reinforcement Learning with Multinomial Logit Function Approximation

### v1

@@ 3 Problem Setup
Assumption 1 is equivalent to the standard assumption in generalized linear model literature (Filippi et al., 2010; Wang et al., 2021) and logistic bandits literature (Faury et al., 2020; Abeille et al., 2021; Zhang and Sugiyama, 2023) to guarantee the Hessian matrix is non-singular.
4 Efficient Algorithm with Online Update
In this part, we design an algorithm that achieves $\mathcal{O}(1)$ computational and storage costs.
Efficiency Concern in Previous Methods.
While Hwang and Oh (2023) made the first step and proposed the UCRL-MNL algorithm with $\widetilde{\mathcal{O}}(\kappa^{-1}dH^{2}\sqrt{K})$ regret, the computational and storage complexity of their algorithm scales linearly with the episode count, which is expensive. Specifically, for the trajectory $\{(s_{k,h},a_{k,h})\}_{h=1}^{H}$ at episode $k$, we define the variable: $y_{k,h}\in\{0,1\}^{N_{k,h}}$ where $y_{k,h}^{s^{\prime}}=\operatorname{\mathds{1}}{(s_{k,h+1}=s^{\prime})}$ for $s^{\prime}\in\mathcal{S}_{k,h}\triangleq\mathcal{S}_{s_{k,h},a_{k,h}}$ and $N_{k,h}=\lvert\mathcal{S}_{k,h}\rvert$. Furthermore, we denote by $p_{k,h}^{s^{\prime}}(\theta)=p_{s_{k,h},a_{k,h}}^{s^{\prime}}(\theta)$. Then $y_{k,h}$ is a sample from the following multinomial distribution:
$y_{k,h}\sim\textnormal{multinomial}(1,[p_{k,h}^{s_{1}}(\theta^{*}),\ldots,p_{k,h}^{s_{N_{k,h}}}(\theta^{*})]).$
where the parameter $1$ indicates that $y_{k,h}$ is a single-trial sample. We define the noise $\epsilon_{k,h}^{s^{\prime}}=y_{k,h}^{s^{\prime}}-p_{k,h}^{s^{\prime}}(\theta_{h}^{*})$. It is clear that $\epsilon_{k,h}\in[-1,1]^{N_{k,h}}$, $\mathbb{E}[\epsilon_{k,h}]=\mathbf{0}$ and $\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\epsilon_{i,h}^{s^{\prime}}=0$.
The UCRL-MNL algorithm estimates the unknown transition parameter $\theta_{h}^{*}$ using the regularized maximum likelihood estimation (MLE). Specifically, based on $y_{k,h}$, the estimator $\theta_{k,h}$ is defined as the solution to the following optimization problem:
@@ 4 Efficient Algorithm with Online Update
In episode $k$, the estimator $\theta_{k,h}$ in (1) is computed using all samples collected in previous episodes, resulting in the computational and storage cost per episode that scales linearly with the episode count, which is expensive in practice.
To address this issue, we introduce a novel algorithm named UCRL-MNL with Online Learning (UCRL-MNL-OL), which achieves both the same regret with constant computational and storage cost. At a high level, our algorithm can be divided into two phases: (i) efficient online estimation, and (ii) efficient optimistic value function construction. We introduce the details in the following.
Efficient Online Estimation.
Instead of using all historical data, our algorithm updates the estimator in an efficient online manner. This is achieved by efficiently exploiting particular properties of the MNL model. The key observation is that the negative log-likelihood function is exponentially concave over the bounded domain, which motivates us to apply a variant of online Newton step (Hazan et al., 2007), in particular, inspired by the works (Hazan et al., 2014; Zhang et al., 2016; Oh and Iyengar, 2021), which focus on logistic bandits and multinomial logit contextual bandits.
First, we define pe-episode loss function $f_{k,h}(\theta)$, gradient $g_{k,h}(\theta)$ and Hessian matrix $H_{k,h}(\theta)$ as
@@
We introduce the following key property of the loss function $f_{k,h}(\theta)$ by a second-order Taylor expansion, whose proof can be found in Lemma 5 in Appendix A.
@@
$H_{k,h}(\theta)\succeq\kappa\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}\triangleq\kappa W_{k,h},\forall\theta\in\Theta.$ (3)
This implies that $f_{k,h}(\theta)$ is exponentially concave over the bounded domain $\Theta$, which enables us to apply the online Newton step to construct an efficient estimator. Specifically, we construct the estimator $\widehat{\theta}_{k+1,h}$ by solving the problem:
@@
where $\widehat{\Sigma}_{k+1,h}=\widehat{\Sigma}_{k,h}+\frac{\kappa}{2}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}$ is the feature covariance matrix. This estimator can be computed efficiently in an online manner without the need to store all historical data. Then, we construct the confidence set by the following lemma.
@@
$\widehat{\mathcal{C}}_{k,h}=\bigg{\{}\theta\in\Theta\mid\lVert\theta-\widehat{\theta}_{k,h}\rVert_{\widehat{\Sigma}_{k,h}}\leq\sqrt{\lambda}B+\sqrt{\frac{8}{\kappa}d\log\left(1+\frac{kUH}{d\lambda\delta}\right)}\triangleq\widehat{\beta}_{k}\bigg{\}}.$
@@
Notably, a similar confidence set is achieved by Hwang and Oh (2023) by using the MLE defined in (1). In contrast, we obtain the same results by using an online estimator which only suffers $\mathcal{O}(1)$ computation and storage cost per round. Besides, we identify a technical issue in their analysis. Specifically, they bound the confidence set using the self-normalized concentration for vector-valued martingales in Lemma 15. However, it is crucial to recognize that the noise is not independent and satisfies $\sum_{s^{\prime}\in\mathcal{S}_{i,h}}\epsilon_{i,h}^{s^{\prime}}=0$ since the learner visits each stage $h$ exactly once per episode. Thus, the noise $\epsilon_{i,h}^{s^{\prime}}$ becomes deterministic and non-zero given the remaining noise $\epsilon_{i,h}^{s^{\prime\prime}}$ for $s^{\prime\prime}\neq s^{\prime}$. This contravenes the zero-mean sub-Gaussian condition in Lemma 15. We observe similar oversights also appear in the works on multinomial logit contextual bandits (Oh and Iyengar, 2019, 2021; Agrawal et al., 2023). To the best of our knowledge, this issue has not been explicitly identified in previous studies. We note that this issue can be resolved with only slight modifications in constant factors by a new self-normalized concentration with dependent noises in Lemma 1 of Li et al. (2024), a simplified version of Theorem C.6 in Périvier and Goyal (2022).
@@
7:        Compute $\widehat{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\langle g_{k,h}(\widehat{\theta}_{k,h}),\theta-\widehat{\theta}_{k,h}\rangle+\frac{1}{2}\lVert\theta-\widehat{\theta}_{k,h}\rVert_{\widehat{\Sigma}_{k+1,h}}^{2}$ as (4).
@@
Efficient Optimistic Value Function Construction.
Given the confidence set $\widehat{\mathcal{C}}_{k,h}$, the most direct way is to adopt the principle of “optimism in the face of uncertainty” and construct the optimistic value function as the maximum expected reward over the confidence set. Specifically, define the value function $\widehat{Q}_{k,h}(s,a)$ and $\widehat{V}_{k,h}(s)$ as
@@
However, this construction is not efficient as it requires solving a maximization problem over the confidence set. Actually, the algorithm only requires the estimated action-value function to be an upper bound for the true value functions. Thus, we can use a closed-form confidence bound instead of computing the maximal value over the confidence set. To this end, we present the following lemma, which enables us to construct the optimistic value function efficiently.
@@
Suppose Lemma 1 holds. For any function $V:\mathcal{S}\to[0,H]$ and $(h,s,a)\in[H]\times\mathcal{S}\times\mathcal{A}$, with probability at least $1-\delta$, it holds that
@@
At state $s_{k,h}$, we chooses $a_{k,h}=\operatorname*{arg\,max}_{a\in\mathcal{A}}\widehat{Q}_{k,h}(s_{k,h},a)$. Algorithm 1 presents the detailed procedure. We show that our algorithm enjoys the following regret guarantee.
@@
Suppose Assumption 1 holds. For any $\delta\in(0,1)$, set $\widehat{\beta}_{k}$ as in Lemma 1 and $\lambda=1$, with probability at least $1-\delta$, Algorithm 1 ensures the following guarantee
${\operatorname*{Reg}}(K)\leq\widetilde{\mathcal{O}}(\kappa^{-1}dH^{2}\sqrt{K}).$
@@
Our algorithm attains the same regret of Hwang and Oh (2023), yet for the more general inhomogeneous MDPs. More importantly, our algorithm only requires $\mathcal{O}(1)$ computational and storage costs, matching the computational efficiency of MDPs with linear function approximation (Ayoub et al., 2020; Jin et al., 2020; Zhou et al., 2021; He et al., 2023).
5 Improved Algorithm by Local Learning
While the UCRL-MNL-OL algorithm enjoys favorable computational efficiency, it suffers from a dependence on $\kappa^{-1}$ in the regret. By Assumption 1, the quantity $\kappa$ may be exponentially small, exhibiting an exponential dependence on the radius of the parameter set and linear dependence on the number of reachable states, i.e., $\kappa^{-1}=\mathcal{O}((U\exp(B))^{2})$. Thus, there is a significant gap between MNL function approximation and linear function approximation. To bridge this gap, we propose new algorithms that leverage local information to enhance the dependence on $\kappa$ significantly.
@@ 5 Improved Algorithm by Local Learning
where the first term is the overestimate of the value function and the second term is the martingale difference sequence due to the stochastic transition. The second term can be upper-bounded by the Azuma-Hoeffding inequality which is independent of $\kappa$. Thus, the $\kappa$-dependence comes from the first term. By the construction of the confidence set in Lemma 1, the dependence on $\widehat{\beta}_{k}$ in the first term leads to a $\kappa^{-1/2}$ dependence. By the configuration of the covariance matrix $\widehat{\Sigma}_{k,h}$ in (4), the dependence on $\widehat{\Sigma}_{k,h}$ also leads to a $\kappa^{-1/2}$ dependence. Thus, we need to design a local learning algorithm that can adaptively use local information to enhance the dependence on these two terms.
@@ 5.1 Improved Online Estimation
In Section 4, we show the Hessian matrix $H_{k,h}$ is lower bounded by a positive definite matrix $\kappa\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}$ uniformly over the parameter domain in (3). This lower bound is not tight as the Hessian matrix can be much larger at certain regions rather than the global minimum. This motivates us to design a local learning algorithm that can adaptively use local information.
Inspired by recent advances in multinomial logistic bandit (Zhang and Sugiyama, 2023) and multinomial logit contextual bandit (Lee and Oh, 2024), we run an online algorithm to estimate the parameter $\theta_{h}^{*}$ similar to (4). Differently, we use the local Hessian matrix ~$\widetilde{\mathcal{H}}_{k,h}$ to update the estimation instead of the global lower bound. Specifically, we estimate parameter $\widetilde{\theta}_{k,h}$ as follows:
@@
where $\eta$ is the step size, $\mathcal{H}_{k,h}=\sum_{i=1}^{k-1}H_{i,h}(\widetilde{\theta}_{i+1,h})+\lambda I_{d}$ and ~$\widetilde{\mathcal{H}}_{k,h}=\mathcal{H}_{k,h}+\eta H_{k,h}(\widetilde{\theta}_{k,h})$. Note both $\mathcal{H}_{k,h}$ and ~$\widetilde{\mathcal{H}}_{k,h}$ can be updated incrementally and the optimization problem in (8) can be solved efficiently. This step can be solved by the following two-step update.
$\theta^{\prime}_{k+1,h}=\theta_{k,h}-\eta\mathcal{H}_{k,h}^{-1}g_{k,h}(\theta_{k,h}),\quad\widetilde{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\lVert\theta-\theta^{\prime}_{k+1,h}\rVert_{\widetilde{\mathcal{H}}_{k,h}}.$ (9)
@@
$\widehat{\mathcal{C}}_{k,h}=\big{\{}\theta\in\Theta\mid\lVert\theta-\widetilde{\theta}_{k,h}\rVert_{\mathcal{H}_{k,h}}\leq\widetilde{\beta}_{k}\big{\}}$
with $\widetilde{\beta}_{k}=\mathcal{O}(\sqrt{d}\log U\log(KH/\delta))$. Then, $\Pr[\theta_{h}^{*}\in\widehat{\mathcal{C}}_{k,h}]\geq 1-\delta,\forall k\in[K],h\in[H]$.
Based on Lemma 3, we can construct the optimistic value function similar to (6) as
$\bar{Q}_{k,h}(s,a)=\bigg{[}r_{h}(s,a)+\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widetilde{\theta}_{k,h})\bar{V}_{k,h+1}(s^{\prime})+H\widetilde{\beta}_{k}\max_{s^{\prime}\in\mathcal{S}_{h,s,a}}\|\phi_{s,a}^{s^{\prime}}\|_{\mathcal{H}_{k,h}^{-1}}\bigg{]}_{[0,H]}.$ (10)
The detailed algorithm is presented in UCRL-MNL-LL in Algorithm 2. We show the regret guarantee of UCRL-MNL-LL by the following theorem.
Theorem 2.
Suppose Assumption 1 holds. For any $\delta\in(0,1)$, set $\eta=\frac{1}{2}\log(1+U)+(B+1)$, $\lambda=84\sqrt{2}\eta(B+d)$, and $\widetilde{\beta}_{k}=\mathcal{O}(\sqrt{d}\log U\log(KH/\delta))$, with probability at least $1-\delta$, UCRL-MNL-LL algorithm ensures the following regret guarantee
${\operatorname*{Reg}}(K)\leq\widetilde{\mathcal{O}}(\kappa^{-1/2}dH^{2}\sqrt{K}).$
Compared with the result in Theorem 1, Theorem 2 enhances the dependence on $\kappa$ from $\kappa^{-1}$ to $\kappa^{-1/2}$. This is achieved by utilizing local information of the Hessian matrix.
Algorithm 2 UCRL-MNL-LL &\&& UCRL-MNL-LL+
@@
3:     UCRL-MNL-LL: Compute $\bar{Q}_{k,h}(\cdot,\cdot)$ in a backward way as in (10).
4:     UCRL-MNL-LL+: Compute $\widetilde{Q}_{k,h}(\cdot,\cdot)$ in a backward way as in (12).
5:     for $h=1,\ldots,H$ do
6:        UCRL-MNL-LL: Observe $s_{k,h}$, select $a_{k,h}=\operatorname*{arg\,max}_{a\in\mathcal{A}}\bar{Q}_{k,h}(s_{k,h},a)$.
7:        UCRL-MNL-LL+: Observe $s_{k,h}$, select $a_{k,h}=\operatorname*{arg\,max}_{a\in\mathcal{A}}\widetilde{Q}_{k,h}(s_{k,h},a)$.
8:        Update ~$\widetilde{\mathcal{H}}_{k,h}=\mathcal{H}_{k,h}+\eta H_{k,h}(\widetilde{\theta}_{k,h})$.
9:        Compute $\widetilde{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\langle g_{k,h}(\widetilde{\theta}_{k,h}),\theta-\widetilde{\theta}_{k,h}\rangle+\frac{1}{2\eta}\lVert\theta-\widetilde{\theta}_{k,h}\rVert_{\widetilde{\mathcal{H}}_{k,h}}$ as (8).
10:        Update $\mathcal{H}_{k,h}=\mathcal{H}_{k,h}+H_{k,h}(\widetilde{\theta}_{k+1,h})$.
11:     end for
12:  end for

### v2

@@ 3 Problem Setup
Assumption 1 is similar to the standard assumption in generalized linear model literature (Filippi et al., 2010; Wang et al., 2021) and logistic bandits literature (Faury et al., 2020; Abeille et al., 2021; Zhang and Sugiyama, 2023) to guarantee the Hessian matrix is non-singular in Property 1.
4 Computationally Efficient Algorithm by Online Learning
In this part, we design an algorithm that achieves $\mathcal{O}(1)$ computational and storage costs per episode.
Since the transition parameter $\theta_{h}^{*}$ is unknown, we need to estimate it using the historical data. At episode $k$, we collect a trajectory $\{(s_{k,h},a_{k,h})\}_{h=1}^{H}$, then define the variable: $y_{k,h}\in\{0,1\}^{N_{k,h}}$ where $y_{k,h}^{s^{\prime}}=\operatorname{\mathds{1}}{(s_{k,h+1}=s^{\prime})}$ for $s^{\prime}\in\mathcal{S}_{k,h}\triangleq\mathcal{S}_{s_{k,h},a_{k,h}}$ and $N_{k,h}=\lvert\mathcal{S}_{k,h}\rvert$. We denote by $p_{k,h}^{s^{\prime}}(\theta)=p_{s_{k,h},a_{k,h}}^{s^{\prime}}(\theta)$. Then $y_{k,h}$ is a sample from the following multinomial distribution:
$y_{k,h}\sim\textnormal{multinomial}(1,[p_{k,h}^{s_{1}}(\theta^{*}),\ldots,p_{k,h}^{s_{N_{k,h}}}(\theta^{*})]),$
where the parameter $1$ indicates that $y_{k,h}$ is a single-trial sample. Furthermore, we define the noise $\epsilon_{k,h}^{s^{\prime}}=y_{k,h}^{s^{\prime}}-p_{k,h}^{s^{\prime}}(\theta_{h}^{*})$. It is clear that $\epsilon_{k,h}\in[-1,1]^{N_{k,h}}$, $\mathbb{E}[\epsilon_{k,h}]=\mathbf{0}$ and $\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\epsilon_{i,h}^{s^{\prime}}=0$.
Efficiency Concern.  Hwang and Oh (2023) made the first step by proposing an approach using the maximum likelihood estimation (MLE). Specifically, their estimator $\theta_{k,h}$ is defined as
@@ 4 Computationally Efficient Algorithm by Online Learning
Despite favorable theoretical guarantee, the computational and storage cost of this method is expensive. First, in episode $k$, the estimator $\theta_{k,h}$ in (1) is computed using all samples collected in previous episodes, resulting an $\mathcal{O}(k)$ storage cost. Second, to solve this optimization problem, the projected gradient descent (Boyd and Vandenberghe, 2004) are usually applied. However, as discussed in Faury et al. (2022), the optimization of the MLE problem typically require $\mathcal{O}(k\log(1/\epsilon))$ iterations to achieve an $\epsilon$-accurate solution, which is computationally expensive.
To address this issue, we introduce a novel algorithm named UCRL-MNL with Online Learning (UCRL-MNL-OL), which achieves both the same regret with $\mathcal{O}(1)$ computational and storage cost. At a high level, our algorithm can be divided into two phases: (i) efficient online estimation, and (ii) efficient optimistic value function construction. We introduce the details in the following.
Efficient Online Estimation.  Instead of using all historical data, our algorithm updates the estimator in an online manner. Specifically, inspired by the works (Hazan et al., 2014; Zhang et al., 2016) on logistic bandit, we find the negative log-likelihood function is exponentially concave, which motivates us to apply a variant of online Newton step (Zhang et al., 2016; Oh and Iyengar, 2021).
First, we define per-episodic loss function $f_{k,h}(\theta)$, gradient $g_{k,h}(\theta)$ and Hessian matrix $H_{k,h}(\theta)$ as
@@
We introduce the following key property of the loss function $f_{k,h}(\theta)$ by a second-order Taylor expansion and mean value theorem, whose proof can be found in Lemma 5 in Appendix A.
@@
$H_{k,h}(\theta)\succeq\kappa\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}\triangleq\kappa W_{k,h}.$ (3)
Property 1 implies that $f_{k,h}(\theta)$ is exponentially concave, which enables us to apply the online Newton step to construct an efficient estimator. Specifically, we construct $\widehat{\theta}_{k+1,h}$ by solving the problem:
@@
where $\widehat{\Sigma}_{k+1,h}=\widehat{\Sigma}_{k,h}+\frac{\kappa}{2}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}$ is the feature covariance matrix.
We show the online estimator $\widehat{\theta}_{k,h}$ in (4) enjoys computational and storage efficiency simultaneously. As the optimization problem exhibits a standard online mirror descent formulation, it can be solved with a projected gradient step with the following equivalent formulation by
$\theta^{\prime}_{k+1,h}=\widehat{\theta}_{k,h}-\widehat{\Sigma}_{k+1,h}^{-1}g_{k,h}(\widehat{\theta}_{k,h}),\quad\mbox{and}\quad\widehat{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\lVert\theta-\theta^{\prime}_{k+1,h}\rVert_{\widehat{\Sigma}_{k+1,h}}.$
This update enjoys a computational cost of only $\mathcal{O}(d^{3}U)$, independent of episode count $k$ (Zhang and Sugiyama, 2023; Lee and Oh, 2024). As for storage costs, it avoids the need to store all historical data by updating the feature covariance matrix $\widehat{\Sigma}_{k,h}$ incrementally, requiring only $\mathcal{O}(d^{2})$ storage cost. We note that a dependence on $U$ is introduced in the computational cost. However, as we have discussed in Remark 1, $U$ can be much smaller than the state space size $S$, which is acceptable. This dependency is typical for model-based methods that directly learn transitions, as it need to control the estimation error of transitions, which typically involves a total of $U$ elements. Similar dependencies have been observed in the literature (Hwang and Oh, 2023). Additionally, even for the model-free method (Yang and Wang, 2020; Zhou et al., 2021) which learn value functions, a common assumption is the value $\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\phi_{k,h}^{s^{\prime}}V(s^{\prime})$ can be obtained by an Oracle, which depends on $U$ implicitly.
Then, we show the estimator $\widehat{\theta}_{k,h}$ is close to the true parameter $\theta_{h}^{*}$ by the following lemma.
@@
$\widehat{\mathcal{C}}_{k,h}=\bigg{\{}\theta\in\Theta\mid\lVert\theta-\widehat{\theta}_{k,h}\rVert_{\widehat{\Sigma}_{k,h}}\leq\sqrt{\lambda}B+\sqrt{\frac{8}{\kappa}d\log\left(1+\frac{kUH}{d\lambda\delta}\right)}\triangleq\widehat{\beta}_{k}\bigg{\}}.$ (5)
@@
Notably, a similar confidence set is achieved by Hwang and Oh (2023) by using the MLE defined in (1). In contrast, we obtain the same results by using an online estimator which only suffers $\mathcal{O}(1)$ computation and storage cost per round. Besides, we identify a technical issue in their analysis. Specifically, they bound the confidence set using the self-normalized concentration for vector-valued martingales in Lemma 15. However, it is crucial to recognize that the noise is not independent and satisfies $\sum_{s^{\prime}\in\mathcal{S}_{i,h}}\epsilon_{i,h}^{s^{\prime}}=0$ since the learner visits each stage $h$ exactly once per episode. Thus, the noise $\epsilon_{i,h}^{s^{\prime}}$ becomes deterministic and non-zero given the remaining noise $\epsilon_{i,h}^{s^{\prime\prime}}$ for $s^{\prime\prime}\neq s^{\prime}$. This contravenes the zero-mean sub-Gaussian condition in Lemma 15. We observe similar oversights also appear in the works on multinomial logit contextual bandits (Oh and Iyengar, 2019, 2021; Agrawal et al., 2023). To our knowledge, this issue has not been explicitly identified in previous studies. We note that this issue can be resolved by a new self-normalized concentration with dependent noises in Lemma 1 of Li et al. (2024b) with only slight modifications in constant factors.
@@
7:         Compute $\widehat{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\langle g_{k,h}(\widehat{\theta}_{k,h}),\theta-\widehat{\theta}_{k,h}\rangle+\frac{1}{2}\lVert\theta-\widehat{\theta}_{k,h}\rVert_{\widehat{\Sigma}_{k+1,h}}^{2}$.
@@
Efficient Optimistic Value Function Construction.  Given the confidence set $\widehat{\mathcal{C}}_{k,h}$, the most direct way is to adopt the principle of “optimism in the face of uncertainty” and construct the optimistic value function as the maximum expected reward over the confidence set. Specifically, we define the optimistic value function $\widehat{Q}_{k,h}(s,a)$ and $\widehat{V}_{k,h}(s)$ as
@@
However, this construction is not efficient as it requires solving a maximization problem over the confidence set. Actually, the algorithm only requires the estimated action-value function to be an upper bound for the true value functions. Thus, we can use a closed-form confidence bound instead of computing the maximal value over the confidence set. To this end, we present the following lemma that enables us to construct the optimistic value function more efficiently.
@@
Suppose Lemma 1 holds. For any $V:\mathcal{S}\to[0,H]$ and $(h,s,a)\in[H]\times\mathcal{S}\times\mathcal{A}$, it holds
@@
At state $s_{k,h}$, our algorithm chooses $a_{k,h}=\operatorname*{arg\,max}_{a\in\mathcal{A}}\widehat{Q}_{k,h}(s_{k,h},a)$. Algorithm 1 presents the detailed procedure. We show that our algorithm achieves the following regret guarantee.
@@
For any $\delta\in(0,1)$, set $\widehat{\beta}_{k}$ as in Lemma 1 and $\lambda=1$, with probability at least $1-\delta$, Algorithm 1 ensures the following regret guarantee
${\operatorname*{Reg}}(K)\leq\widetilde{\mathcal{O}}\big{(}\kappa^{-1}dH^{2}\sqrt{K}\big{)}.$
@@
Our algorithm attains the same regret of Hwang and Oh (2023), yet for the more general inhomogeneous MDPs. Importantly, our algorithm only requires constant computational and storage costs, matching the computational efficiency of the linear cases (Jin et al., 2020; Zhou et al., 2021).
5 Statistically Improved Algorithm by Local Learning
In this section, we present an enhanced algorithm, UCRL-MNL-LL, that leverages local information to improve the statistical efficiency for MDPs with MNL function approximation.
While the UCRL-MNL-OL algorithm in Section 4 offers favorable computational and storage efficiency, it suffers from a dependence on $\kappa^{-1}$ in the regret. By Assumption 1, the quantity $\kappa$ may be exponentially small in the worst case, exhibiting an exponential dependence on the radius of the parameter set and linear in the number of reachable states, i.e., $\kappa^{-1}=\mathcal{O}((U\exp(B))^{2})$. This creates a significant gap between MNL and linear function approximation.
Recently, the improved dependence on $\kappa$ has been achieved in the logistic bandit literature (Faury et al., 2020; Abeille et al., 2021; Périvier and Goyal, 2022) by the use of generalization of the Bernstein-like tail inequality (Faury et al., 2020) and the self-concordant-like property of the log loss (Bach, 2010). Thus, a natural question then arises: Can we improve the statistical efficiency of MNL function approximation for MDPs? We answer this question affirmatively by proposing an enhanced algorithm that reduces the dependence on $\kappa$ significantly through the use of local information. Though the achievement has been made in the logistic bandit, the extension to MDPs with MNL function approximation is non-trivial and new techniques specific to MDPs are required.
@@ 5 Statistically Improved Algorithm by Local Learning
Here, the first term corresponds to the overestimation of the value function, and the second term represents the martingale difference sequence arising from the stochastic transition dynamics. The second term can be bounded using the Azuma-Hoeffding inequality, which is independent of $\kappa$. For the first term, by the construction of the confidence set in Lemma 1, $\widehat{\beta}_{k}$ represents the width of the confidence set, contributing to a $\kappa^{-1/2}$ dependence. The feature covariance matrix $\widehat{\Sigma}_{k,h}$ defined in (4) reflects the degree of exploration across different states, which also results in a $\kappa^{-1/2}$ dependence.
To mitigate these dependencies, we design a local learning algorithm that: (i) constructs a confidence set independent of $\kappa$; and (ii) builds a $\kappa$-independent feature covariance matrix that effectively captures the exploration degree across different states. We introduce the details in the following.
@@ 5.1 Improved Online Estimation
In Property 1, we show the Hessian matrix $H_{k,h}$ is lower bounded by a positive definite matrix $\kappa\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}$ uniformly over the parameter domain, as in (3). This quantity measures the exploration degree across different states and is used to update the parameter estimation and construct the confidence set. However, the bound is not tight, as the Hessian matrix can be significantly larger in certain regions, away from the global minimum. This observation motivates us to design a local learning algorithm that can adaptively leverage local information for improved guarantee.
Inspired by recent advances in multinomial logistic bandit (MLogB) (Zhang and Sugiyama, 2023) and multinomial logit contextual bandit (MNL) (Lee and Oh, 2024), we run an online mirror descent algorithm to estimate the parameter $\theta_{h}^{*}$. Differently from (4), we use the local Hessian matrix ~$\widetilde{\mathcal{H}}_{k,h}$ to update the estimation instead of the global lower bound. Specifically, we estimate $\widetilde{\theta}_{k,h}$ as follows:
@@
where $\eta$ is the step size, $\mathcal{H}_{k,h}=\sum_{i=1}^{k-1}H_{i,h}(\widetilde{\theta}_{i+1,h})+\lambda I_{d}$ and ~$\widetilde{\mathcal{H}}_{k,h}=\mathcal{H}_{k,h}+\eta H_{k,h}(\widetilde{\theta}_{k,h})$. Note that both $\mathcal{H}_{k,h}$ and ~$\widetilde{\mathcal{H}}_{k,h}$ can be updated incrementally. Similar to the update in (4), the optimization problem in (9) can be efficiently solved using the following two-step update:
$\theta^{\prime}_{k+1,h}=\theta_{k,h}-\eta\widetilde{\mathcal{H}}_{k,h}^{-1}g_{k,h}(\theta_{k,h}),\quad\mbox{and}\quad\widetilde{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\lVert\theta-\theta^{\prime}_{k+1,h}\rVert_{\widetilde{\mathcal{H}}_{k,h}}.$
This two-step update procedure incurs a computational cost of $\mathcal{O}(d^{3}U)$ and a storage cost of $\mathcal{O}(d^{2})$ per episode, both independent of the episode count $k$.
@@
$\widetilde{\mathcal{C}}_{k,h}=\big{\{}\theta\in\Theta\mid\lVert\theta-\widetilde{\theta}_{k,h}\rVert_{\mathcal{H}_{k,h}}\leq\widetilde{\beta}_{k}\big{\}},$
where $\widetilde{\beta}_{k}=\mathcal{O}(\sqrt{d}\log U\log(KH/\delta))$. Then, we have $\Pr[\theta_{h}^{*}\in\widetilde{\mathcal{C}}_{k,h}]\geq 1-\delta,\forall k\in[K],h\in[H]$.
Remark 5.
Zhang and Sugiyama (2023) studied the multiple-parameter MLogB model, where the unknown parameter is a matrix. Consequently, the confidence set in Theorem 3 of their work exhibits a polynomial dependence on the number of possible outcomes, which corresponds to the number of reachable states $U$ in our setting. This dependence is acceptable in the bandit setting while is not suitable for the MDP setting. In contrast, Lee and Oh (2024) focused on the single-parameter MNL model and revisited the self-concordant-like property, demonstrating that the log-loss of the single-parameter MNL model is $3\sqrt{2}$-self-concordant-like (Proposition B.1 in Lee and Oh (2024)). This property is crucial for the improved confidence set in Lemma 3 that is independent of $\kappa$ and $U$.
Algorithm 2 UCRL-MNL-LL
@@
3:     Compute $\widetilde{Q}_{k,h}(\cdot,\cdot)$ in a backward way as in (11).
4:     for $h=1,\ldots,H$ do
5:         Observe state $s_{k,h}$, select action $a_{k,h}=\operatorname*{arg\,max}_{a\in\mathcal{A}}\widetilde{Q}_{k,h}(s_{k,h},a)$.
6:         Update ~$\widetilde{\mathcal{H}}_{k,h}=\mathcal{H}_{k,h}+\eta H_{k,h}(\widetilde{\theta}_{k,h})$.
7:         Compute $\widetilde{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\langle g_{k,h}(\widetilde{\theta}_{k,h}),\theta-\widetilde{\theta}_{k,h}\rangle+\frac{1}{2\eta}\lVert\theta-\widetilde{\theta}_{k,h}\rVert_{\widetilde{\mathcal{H}}_{k,h}}$.
8:         Update $\mathcal{H}_{k,h}=\mathcal{H}_{k,h}+H_{k,h}(\widetilde{\theta}_{k+1,h})$.
9:     end for
10:  end for